OpenStudy (anonymous):

What is 4/5 of 10/13? I'm sorry if this seems too easy, but I am kind of sick and I haven't been able to focus well. The answer should be in fraction form. Thanks!

OpenStudy (anonymous):

I well medal anyone with the correct answer:)

OpenStudy (anonymous):

Answer choices?

OpenStudy (anonymous):

40/65 5/18 8/13 20/65

OpenStudy (jhannybean):

\[\large \frac{4}{5} \cdot \frac{10}{13} = \frac{4}{\cancel{5}} \cdot \frac{\cancel{10}2}{13} = \frac{4 \cdot 2}{13} =~?\]

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

Lets not give direct answers ^_^ please explain thank you

OpenStudy (anonymous):

Thanks to Jhannybean! I like it a lot better when people explain rather than just give out answers ^.^

OpenStudy (anonymous):

Why should I bother as right when I typed the answer and sent it, JhannyBean explained/. @Captain_America_Army Please, leave my responses B. Thank yew.

OpenStudy (jhannybean):

Np.